7) The reaction below is exothermic 2502 (8) + 02 (8) 2503 (8) will result in an increase in the number of moles of SO3 (8) i Just give me the answers
Answer #1


-7) 25029) + O2(g) = 2509 (g) + xuj exothermic heat content A increasing the volume of container means reducing pressure of s- a > Profonation af NH3 giver what .. conseok odd 4 NH, is NHÀ 10> [ok] = 0.0025Mo s issa: POH = -log Cow] lik? = -log (0.00